# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ 2.1 Quadratic action and primordial power spectra with non-Bunch-Davies states
$\langle\hat{\zeta}({\bf k})\hat{\zeta}({\bf k}^{\prime})\rangle=(2\pi)^{3}\delta({\bf k}+{\bf k}^{\prime})\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{\zeta},$ (2.23)
$\langle\hat{\xi}^{(s)}({\bf k})\hat{\xi}^{(s^{\prime})}({\bf k}^{\prime})\rangle=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}+{\bf k}^{\prime})\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h},$ (2.24)
where $\hat{\xi}^{(s)}({\bf k}):=\hat{h}_{ij}({\bf k})e^{(s)*}_{ij}({\bf k})$. Then, the power spectra on superhorizon scales ($-c_{s}k\eta\ll 1$ and $-c_{h}k\eta\ll 1$), can be obtained from the mode functions as
$\mathcal{P}_{\zeta}$ $=\frac{1}{8\pi^{2}}\frac{H^{2}}{\mathcal{F}_{S}c_{s}}|\alpha_{k}-\beta_{k}|^{2},$ (2.25)
$\mathcal{P}^{(s)}_{h}$ $=\frac{1}{\pi^{2}}\frac{H^{2}}{\mathcal{F}_{T}c_{h}}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (2.26)
@@
$n_{s}-1$ $:=\frac{{\rm d}\ln\mathcal{P}_{\zeta}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{S}}{2}-\frac{3}{2}f_{S}+\frac{{\rm d}}{{\rm d}\ln k}\ln|\alpha_{k}-\beta_{k}|^{2},$ (2.27)
$n_{t}$ $:=\frac{{\rm d}\ln\mathcal{P}_{h}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{T}}{2}-\frac{3}{2}f_{T}+\frac{{\rm d}}{{\rm d}\ln k}\ln\sum_{s}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (2.28)
where $\mathcal{P}_{h}=\sum_{s}\mathcal{P}^{(s)}_{h}$. By taking the limit, $\alpha_{k}=1,~{}\beta_{k}=0,~{}\alpha^{(s)}_{k}=1$, and $\beta^{(s)}_{k}=0$, Eqs. (2.25)–(2.28) reproduce the results in the slow-roll inflation with the Bunch-Davies state [82]. In the main text, we hereafter set $\alpha_{k}=1$ and $\beta_{k}=0$ (i.e., assume the Bunch-Davies scalar modes) for simplicity, and we solely study the impact of the non-Bunch-Davies tensor modes on the SGWB and its anisotropies. (See Ref. [85] for the impacts of primordial gravitational waves induced by non-Bunch-Davies scalar modes on the SGWB.)
@@ 2.2 On the scale-dependence of the tensor power spectrum with non-Bunch-Davies state
The last term in Eq. (2.28) is peculiar to the excited tensor modes, and $k$ dependence of the Bogoliubov coefficients can yield a variety of non-standard predictions, e.g., a highly blue-tilted spectrum [86, 84]. Such a blue-tilted spectrum can potentially be tested by PTA experiments such as, e.g., NANOGrav, as has been studied in Ref. [84] where a power-law $k$ dependence of the power spectrum was assumed. On the other hand, the nearly-scale-invariant spectrum also has a chance to be tested by planned space interferometers such as DECIGO/BBO. Thus, here, we introduce phenomenological models for the Bogoliubov coefficients to realize sufficient amplitude of the SGWB to detect in both frequency bands. (See, e.g., Refs. [87, 88, 89, 90, 91] where impacts of non-Bunch-Davies states on scalar/tensor power spectra have been studied with different Bogoliubov coefficients.)
@@ 2.2.1 Peaked spectrum
$\alpha^{(s)}_{k}$ $=1+i{\rm Im}[\alpha^{(s)}_{k}],$ (2.29)
$\beta^{(s)}_{k}$ $={\rm Re}[\beta^{(s)}_{k}].$ (2.30)
From the normalization condition Eq. (2.22), one can derive
$({\rm Im}[\alpha^{(s)}_{k}])^{2}=({\rm Re}[\beta^{(s)}_{k}])^{2}.$ (2.31)
@@
${\rm Re}[\beta^{(s)}_{k}]=A\times\exp[-\frac{(\ln{k/k_{*}})^{2}}{\Delta^{2}}\biggr{]},$ (2.32)
where $A$ and $\Delta$ are dimensionless parameters that control how large and how sharp the peak is. For $A=-3000$ and $\Delta=0.45$, we can realize a blue-tilted tensor spectrum with $n_{t}\simeq 2$ for $k\lesssim k_{*}$, and an $\mathcal{O}(10^{7})$ enhancement in the power spectrum at a reference scale $k=k_{*}$. Ref. [84] mentioned that the nHz SGWB detected by recent PTA experiments [92, 93, 94, 4, 95, 96, 97, 98, 99, 100] can be explained by the single-field inflation with the non-Bunch-Davies vacuum state with choosing $k_{*}\sim 10^{8}{\rm Mpc}^{-1}$.
2.2.2 Nearly-scale-invariant spectrum
For a nearly-scale-invariant case, we employ the following ansatz: [^2] Here, we assumed a pure imaginary $\beta^{(s)}_{k}$. This is because the scale-invariant case with a pure real $\beta^{(s)}_{k}$ yields a leading-order contribution of $\mathcal{O}(B^{2})$ in Eqs. (4.48) and (4.49). As will be explained later, one has $B\ll 1$ due to the backreaction constraint. Therefore, we need a pure imaginary $\beta^{(s)}_{k}$ that leads to the leading-order contribution of $\mathcal{O}(B)$ in Eqs. (4.48) and (4.49).
$\alpha^{(s)}_{k}$ $=1+iB,$ (2.33)
$\beta^{(s)}_{k}$ $=iB,$ (2.34)
where $B$ is real and constant. In this case, the last term in Eq. (2.28) is zero, and the spectral index of the tensor power spectrum is given by the slow-roll parameters.
@@ 3 Non-Bunch-Davies bispectrum
As we will see later in section 4, the SGWB anisotropies on large scales are sensitive to the primordial non-Gaussianity which gives the coupling between short- and long-wavelength modes (see, e.g., Ref. [29, 32]). At the leading order, such a coupling can be characterized by the primordial bispectrum in the squeezed limit. Although the tensor auto-bispectrum in the squeezed limit can be a source of the anisotropies of primordial GWB, due to the tight observational constraint on the primordial tensor modes on CMB scales the contribution from the tensor auto-bispectrum might be smaller than that from the STT cross-bispectrum. Then, here, we focus on the primordial STT bispectrum in the squeezed limit, that is, the coupling between long-wavelength scalar modes and short-wavelength tensor modes.
3.1 Trans-Planckian problem
Before formulating the STT bispectrum, we mention the Trans-Planckian issue [79] in the inflation models with the non-Bunch Davies state. In general, the physical momenta of short-wavelength modes, e.g., GWs targeted by PTA such as NANOGrav or space interferometer such as DECIGO and BBO, can exceed the Planck scale (i.e., $k_{\rm phys}:=k/a\geq M_{\rm pl}$) when both short- and long-wavelength modes are on subhorizon scales at the same time. The condition that the long-wavelength modes are on subhorizon scales at $\eta=\eta_{0}$ is expressed as
$1<-k_{L}\eta_{0}\,\left(\,\ll-k_{S}\eta_{0}\,\right),$ (3.1)
@@ 3.1 Trans-Planckian problem
where $k_{S}$ and $k_{L}$ denote the comoving wavenumber of the short- and long-wavelength modes, respectively. In the quasi-de Sitter background (i.e., $a\simeq-(H\eta)^{-1}$), Eq. (3.1) turns into
$\frac{k_{{\rm phys},L}}{M_{\rm pl}}>\frac{H_{\rm inf}}{M_{\rm pl}},$ (3.2)
and for the short-wavelength modes
$\frac{k_{{\rm phys},S}}{M_{\rm pl}}>\frac{H_{\rm inf}}{M_{\rm pl}}\times\frac{k_{S}}{k_{L}}.$ (3.3)
Assuming $\mathcal{O}(\mathcal{G}_{T},\mathcal{F}_{T})=M_{\rm Pl}^{2}$ and $|\alpha^{(s)}_{k_{\rm CMB}}-\beta^{(s)}_{k_{\rm CMB}}|^{2}\sim 1$, Eq. (2.26) gives $\mathcal{P}_{h}^{1/2}|_{k_{\rm CMB}}\sim H_{\rm inf}/M_{\rm pl}$, and then we have
$10^{-6}\left(\frac{r|_{k_{\rm CMB}}}{10^{-3}}\right)^{1/2}\left(\frac{\mathcal{P}_{\zeta}|_{k_{\rm CMB}}}{2\times 10^{-9}}\right)^{1/2}\left(\frac{k_{S}}{k_{L}}\right)<\frac{k_{{\rm phys},S}}{M_{\rm pl}},$ (3.4)
where $r:=\mathcal{P}_{h}/\mathcal{P}_{\zeta}$ is the tensor-to-scalar ratio. It is expected that future CMB B-mode polarization experiments can detect the inflationary SGWB with $r|_{k_{\rm CMB}}\sim 10^{-3}$, and then the above inequality means that the detection of the inflationary SGWB would suffer from the Trans-Planckian problem of the short-wavelength modes at $\eta=\eta_{0}$ for $k_{S}>10^{4}~{}{\rm Mpc}^{-1}\times(k_{L}/(10^{-2}~{}{\rm Mpc}^{-1}))$. Primordial non-Gaussianities from inflation with the Bunch-Davies state are insensitive to subhorizon modes since those are mainly generated after their horizon-cross scales, while those with the non-Bunch-Davies state are not. This is because the excited modes on subhorizon scales can generate the non-Gaussianities [39] and then the above Trans-Planckian issue becomes serious.
@@
In principle, to avoid this problem one can consider a non-standard model of inflation with $\mathcal{P}_{\zeta}|_{k_{\rm CMB}}\sim 10^{-9}$ and $r\ll 10^{-3}$ by choosing model parameters, but we do not consider such models. Instead, we relax the starting point of the above argument Eq. (3.1) in such a way that the tensor modes got excited from the Bunch-Davies state at $\eta=\eta_{0}$ when the long-wavelength modes of $k_{L}=10^{-2}{\rm Mpc}^{-1}$ were already on superhorizon scales, while the short-wavelength modes were still on subhorizon scales:
$-k_{S}\eta_{0}>1\gg-k_{L}\eta_{0},$ (3.5)
and hence we obtain
$\frac{k_{{\rm phys},S}}{M_{\rm pl}}>r^{1/2}\mathcal{P}^{1/2}_{\zeta}|_{k_{\rm CMB}}\gg\frac{k_{{\rm phys},L}}{M_{\rm pl}}.$ (3.6)
@@
By choosing $\eta_{0}$ appropriately, the Trans-Planckian problem can be avoided even for $k_{S}\gg 10\,{\rm Mpc}^{-1}$, e.g., the modes corresponding to the ones targeted by NANOGrav ($10^{6}\,\rm Mpc^{-1}$) and DECIGO/BBO ($10^{13}\,\rm Mpc^{-1}$). It has been extensively discussed that the non-Bunch-Davies tensor modes can enhance primordial bispectra compared to those in the case of the Bunch-Davies state. In particular, the enhancements of the so-called squeezed bispectra are in proportion to powers of $k_{S}/k_{L}$ (see, e.g., Ref. [51]). However, we will show that this enhancement appears in the case with Eq. (3.1), while it does not with Eq. (3.5).
3.2 Scalar-tensor-tensor bispectrum
The scalar-tensor-tensor bispectrum is defined by
$\langle\hat{\zeta}({\bf k}_{1})\hat{\xi}^{(s_{2})}({\bf k}_{2})\hat{\xi}^{(s_{3})}({\bf k}_{3})\rangle=(2\pi)^{3}\delta({\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3})\mathcal{B}_{shh}^{s_{2}s_{3}}.$ (3.7)

### v3

@@ 2.1 Quadratic action and primordial power spectra with non-Bunch-Davies states
$\langle\hat{\zeta}({\bf k})\hat{\zeta}({\bf k}^{\prime})\rangle$ $=(2\pi)^{3}\delta({\bf k}+{\bf k}^{\prime})\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{\zeta},$ (23)
$\langle\hat{\xi}^{(s)}({\bf k})\hat{\xi}^{(s^{\prime})}({\bf k}^{\prime})\rangle$ $=(2\pi)^{3}\delta_{ss^{\prime}}\delta({\bf k}+{\bf k}^{\prime})\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h}$, (24)
where $\hat{\xi}^{(s)}({\bf k}):=\hat{h}_{ij}({\bf k})e^{(s)*}_{ij}({\bf k})$. Then, the power spectra on superhorizon scales ($c_{s}k\eta\ll 1$ and $c_{h}k\eta\ll 1$), can be obtained from the mode functions as
$\mathcal{P}_{\zeta}$ $=\frac{1}{8\pi^{2}}\frac{H^{2}}{\mathcal{F}_{S}c_{s}}|\alpha_{k}-\beta_{k}|^{2},$ (25)
$\mathcal{P}^{(s)}_{h}$ $=\frac{1}{\pi^{2}}\frac{H^{2}}{\mathcal{F}_{T}c_{h}}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (26)
@@
$n_{s}-1$ $:=\frac{{\rm d}\ln\mathcal{P}_{\zeta}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{S}}{2}-\frac{3}{2}f_{S}+\frac{{\rm d}}{{\rm d}\ln k}\ln|\alpha_{k}-\beta_{k}|^{2},$ (27)
$n_{t}$ $:=\frac{{\rm d}\ln\mathcal{P}_{h}}{{\rm d}\ln k}\simeq-2\epsilon+\frac{g_{T}}{2}-\frac{3}{2}f_{T}+\frac{{\rm d}}{{\rm d}\ln k}\ln\sum_{s}|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2},$ (28)
where $\mathcal{P}_{h}=\sum_{s}\mathcal{P}^{(s)}_{h}$. By taking the limit, $\alpha_{k}=1,~\beta_{k}=0,~\alpha^{(s)}_{k}=1$, and $\beta^{(s)}_{k}=0$, Eqs. (25)–(28) reproduce the results in the slow-roll inflation with the Bunch-Davies state Kobayashi:2011nu . In the main text, we hereafter set $\alpha_{k}=1$ and $\beta_{k}=0$ (i.e. assume the Bunch-Davies scalar modes) for simplicity, and we solely study the impact of the non-Bunch-Davies tensor modes on the SGWB and its anisotropies. (See Ref. Fumagalli:2021mpc for the impacts of primordial GWs induced by non-Bunch-Davies scalar modes on the SGWB.)
@@ 2.2 On the scale-dependence of the tensor power spectrum with non-Bunch-Davies state
The last term in Eq. (28) is peculiar to the excited tensor modes, and $k$-dependence of the Bogoliubov coefficients can yield a variety of non-standard predictions, e.g. a highly blue-tilted spectrum Ashoorioon:2014nta ; Choudhury:2023kam . Such a blue-tilted spectrum can potentially be tested by PTA experiments such as NANOGrav, as has been studied in Ref. Choudhury:2023kam where a power-law $k$-dependence of the power spectrum was assumed. On the other hand, the nearly scale-invariant spectrum also has a chance to be tested by planned space interferometers such as DECIGO/BBO. Thus, here, we introduce phenomenological models for the Bogoliubov coefficients to realize sufficient amplitude of the SGWB for detection in both frequency bands. (See, e.g. Refs. Danielsson:2002kx ; Goldstein:2002fc ; Alberghi:2003am ; Broy:2016zik ; Cielo:2022vmo where impacts of non-Bunch-Davies states on scalar/tensor power spectra have been studied with different Bogoliubov coefficients.)
@@ 2.2.1 Peaked spectrum
$\alpha^{(s)}_{k}$ $=1+i{\rm Im}[\alpha^{(s)}_{k}$], (29)
$\beta^{(s)}_{k}$ $={\rm Re}[\beta^{(s)}_{k}$]. (30)
From the normalization condition Eq. (22), one can derive
$({\rm Im}[\alpha^{(s)}_{k}])^{2}=({\rm Re}[\beta^{(s)}_{k}])^{2}.$ (31)
@@
[…]
where $A$ and $\Delta$ are dimensionless parameters that control how large and how sharp the peak is. Here, the above ansatz covers various types of the GW spectrum such as an almost flat spectrum and a weaker or broader peaked one. However, among those, one needs to take $A$ to be much smaller than unity to get the flat spectrum, which suppresses the enhancement of anisotropies as will be shown later. Therefore, as an exception, we consider a different ansatz leading to the flat spectrum that can enhance the anisotropies in the following subsection.
For A=3000A=-3000 and $\Delta=0.45$, we can realize a blue-tilted tensor spectrum with $n_{t}\simeq 2$ for $k\lesssim k_{*}$, and an $\mathcal{O}(10^{7})$ enhancement in the power spectrum at a reference scale $k=k_{*}$. Ref. Choudhury:2023kam mentioned that the nHz SGWB detected by recent PTA experiments NANOGrav:2020bcs ; Goncharov:2021oub ; EPTA:2021crs ; EPTA:2023fyk ; EPTA:2023sfo ; EPTA:2023akd ; EPTA:2023gyr ; EPTA:2023xxk ; EuropeanPulsarTimingArray:2023egv ; Antoniadis:2022pcn can be explained by the single-field inflation with the non-Bunch-Davies vacuum state if choosing $k_{*}\sim 10^{8}{\rm Mpc}^{-1}$. Here, one might consider different values of $\Delta$ to realize a different sharpness of the GW spectrum with enhanced anisotropies. Note that, as we will see later, we have to be careful about the backreaction from the excited states on the inflationary background dynamics (see, e.g. Ref. Akama:2020jko ). For $\Delta\gtrsim 1$ to realize a wider peak, more modes contribute to the backreaction, and thus we will get a more stringent constraint on the non-Bunch-Davies states compared to the above choice. In fact, as will be shown later, one cannot arbitrarily increase the value of $\eta_{0}$, which is introduced as an excitation time of the Bunch-Davies tensor modes, due to Trans-Planckian and strong coupling issues, and the maximum possible value of $k\eta_{0}$ for $\Delta=0.45$ is almost the same as that determined by requiring the absence of those issues. Therefore, to discuss the highest detectability of anisotropies by the PTA experiments, we do not consider the other values.
2.2.2 Nearly scale-invariant spectrum
As we mentioned before, DECIGO can potentially test a nearly scale-invariant spectrum. Thus, as the minimum setup for the case of DECIGO, we consider a nearly flat spectrum by employing the following ansatz: [^2] Here, we assumed a pure imaginary $\beta^{(s)}_{k}$. This is because the scale-invariant case with a pure real $\beta^{(s)}_{k}$ yields a leading-order contribution of $\mathcal{O}(B^{2})$ in Eqs. (130) and (131). As will be explained later, one has $B\ll 1$ due to the backreaction constraint. Therefore, we need a pure imaginary $\beta^{(s)}_{k}$ that leads to the leading-order contribution of $\mathcal{O}(B)$ in Eqs. (130) and (131).
$\alpha^{(s)}_{k}$ $=1+iB,$ (33)
$\beta^{(s)}_{k}$ $=iB,$ (34)
where $B$ is real and constant. In this case, the last term in Eq. (28) is zero, and the spectral index of the tensor power spectrum is given by the slow-roll parameters.
@@ 3 Non-Bunch-Davies bispectrum
As we will see in section 4, the SGWB anisotropies on large scales are sensitive to the primordial non-Gaussianity which gives the coupling between short- and long-wavelength modes (see, e.g. Refs. Dimastrogiovanni:2021mfs ; Orlando:2022rih ). At the leading order, such a coupling can be characterized by the primordial bispectrum in the squeezed limit. Although the tensor auto-bispectrum in the squeezed limit can be a source of the anisotropies of primordial GWB, due to the tight observational constraint on the primordial tensor modes on CMB scales the contribution from the tensor auto-bispectrum might be smaller than that from the STT cross-bispectrum. Then, here, we focus on the primordial STT bispectrum in the squeezed limit, i.e. the coupling between long-wavelength scalar modes and short-wavelength tensor modes.
3.1 Trans-Planckian and strong coupling problems
Before formulating the STT bispectrum, we mention the Trans-Planckian issue Martin:2000xs and a strong coupling issue in the inflation models with the non-Bunch-Davies state.
In general, in the inflationary era, the physical momenta of each mode could exceed the Planck scale. In particular, this issue may become more serious in the situation considered here, where the large hierarchy between the long- and short-wavelength modes exists. In addition, Ref. Kunimitsu:2015faa pointed out that in the Horndeski theory the strong coupling scale should generally appear below the Planck scale and it is estimated as
$\Lambda\simeq(\sqrt{\epsilon}M_{\rm Pl}H^{2})^{1/3}\ll M_{\rm Pl},$ (35)
where $\mathcal{G}_{S}/M_{\rm Pl}^{2},\mathcal{F}_{S}/M_{\rm Pl}^{2}\sim\epsilon\ll 1$ and $H/M_{\rm Pl}\ll 1$ are assumed. It should be noted that, depending on the choice of the arbitrary functions in Horndeski’s action (1), a higher cutoff scale can be obtained. Hereafter, we employ the lowest strong coupling scale given by Eq. (35), which gives the most conservative condition on the excitation time of the tensor modes, $\eta_{0}$.
Thus, we need to require that the physical momenta, in particular, of the short-wavelength modes do not exceed these scales. We discuss this issue separately in two situations. First, let us consider the case where both short- and long-wavelength modes are on subhorizon scales at a certain time during inflation. The condition that the long-wavelength modes are on subhorizon scales at $\eta=\eta_{0}$ is expressed as
$1<-k_{L}\eta_{0}\,\left(\,\ll-k_{S}\eta_{0}\,\right),$ (36)
@@ 3.1 Trans-Planckian and strong coupling problems
where $k_{S}$ and $k_{L}$ denote the comoving wavenumber of the short- and long-wavelength modes, respectively. In the quasi-de Sitter background (i.e. $a\simeq-(H\eta)^{-1}$), Eq. (36) turns into
$k_{{\rm phys},L}>H_{\rm inf},~k_{{\rm phys},S}>H_{\rm inf}\times\frac{k_{S}}{k_{L}},$~ (37)
at $\eta=\eta_{0}$. As for the strong coupling issue, it should be required that
$1>\frac{k_{{\rm phys},S}}{\Lambda}~,$ (38)
and hence we can obtain the requirement for the hierarchy between the short- and long-wavelength modes as
$\frac{\Lambda}{H_{\rm inf}}>\frac{k_{S}}{k_{L}}~.$ (39)
By using the expression for the cutoff scale in generalized G-inflation (35), we have
$\left(\frac{\sqrt{\epsilon}M_{\rm Pl}}{H_{\rm inf}}\right)^{1/3}>\frac{k_{S}}{k_{L}}~,$ (40)
and this inequality can be estimated as
$O(10)\times\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right)^{-1/6}>\frac{k_{S}}{k_{\rm CMB}}~,$ (41)
@@
where we have used the expression for the power spectrum of the curvature perturbations (25) and assumed $\mathcal{G}_{S}/M_{\rm Pl}^{2},\mathcal{F}_{S}/M_{\rm Pl}^{2}\sim\epsilon$ and $|\alpha_{k_{\rm CMB}}-\beta_{k_{\rm CMB}}|^{2}\sim 1$. Here, we take $k_{L}$ as the CMB scale. This inequality means that the detection of the inflationary SGWB would suffer from the strong coupling problem of the short-wavelength modes at $\eta=\eta_{0}$ for $k_{S}>O(10)\times k_{\rm CMB}$. To avoid this problem, let us consider the case that relaxes the starting condition of the above argument based on Eq. (36), in such a way that the tensor modes with $k_{S}$ got excited from the Bunch-Davies state at $\eta=\eta_{0}$ when the long-wavelength modes of $k_{L}$ (typically, CMB scale) were already on superhorizon scales, while the short-wavelength modes were still on subhorizon scales:
$-k_{S}\eta_{0}>1\gg-k_{L}\eta_{0}.$ (42)
The strong coupling problem requires
$\frac{\Lambda}{H_{\rm inf}}>-k_{S}\eta_{0},$ (43)
at $\eta=\eta_{0}$, where we have used $a\simeq-(H\eta)^{-1}$. By employing Eq. (35), we have
$\left(\frac{\sqrt{\epsilon}M_{\rm Pl}}{H_{\rm inf}}\right)^{1/3}>-k_{S}\eta_{0}~,$ (44)
and this inequality can be estimated as
$O(10)\times\left(\frac{\mathcal{P}_{\zeta}}{2\times 10^{-9}}\right)^{-1/6}>-k_{S}\eta_{0}.$ (45)
Here, we introduce $\eta_{0}$ as the time when the tensor modes got excited from the Bunch-Davies vacuum state. Then, if we properly choose $\eta_{0}$ so that $1<-k_{S}\eta_{0}\leq O(10)$ is satisfied, the aforementioned issues can be absent even in the case where the short-wavelength tensor modes are on the subhorizon scales at $\eta=\eta_{0}$.
@@
It has been extensively discussed that the non-Bunch-Davies tensor modes can enhance primordial bispectra compared to those in the case of the Bunch-Davies state. In particular, the enhancements of the so-called squeezed bispectra are in proportion to powers of $k_{S}/k_{L}$ (see, e.g. Ref. Ganc:2012ae ). However, based on the above discussion, we would like to stress that this enhancement appears in the case with Eq. (36), whereas it does not with Eq. (42).
3.2 STT bispectrum
The STT bispectrum is defined by
$\langle\hat{\zeta}({\bf k}_{1})\hat{\xi}^{(s_{2})}({\bf k}_{2})\hat{\xi}^{(s_{3})}({\bf k}_{3})\rangle=(2\pi)^{3}\delta({\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3})\mathcal{B}_{shh}^{s_{2}s_{3}}.$ (46)